# Reinforcement Learning Outperforms Supervised Fine-Tuning: A Case Study on Audio Question Answering

### v1

@@ Abstract
Recently, reinforcement learning (RL) has been shown to greatly enhance the reasoning capabilities of large language models (LLMs), and RL-based approaches have been progressively applied to visual multimodal tasks. However, the audio modality has largely been overlooked in these developments. Thus, we conduct a series of RL explorations in audio understanding and reasoning, specifically focusing on the audio question answering (AQA) task. We leverage the group relative policy optimization (GRPO) algorithm to Qwen2-Audio-7B-Instruct, and our experiments demonstrated state-of-the-art performance on the MMAU Test-mini benchmark, achieving an accuracy rate of 64.5%. The main findings in this technical report are as follows: 1) The GRPO algorithm can be effectively applied to large audio language models (LALMs), even when the model has only 8.2B parameters; 2) With only 38k post-training samples, RL significantly outperforms supervised fine-tuning (SFT), indicating that RL-based approaches can be effective without large datasets; 3) The explicit reasoning process has not shown significant benefits for AQA tasks, and how to efficiently utilize deep thinking remains an open question for further research; 4) LALMs still lag far behind humans auditory-language reasoning, suggesting that the RL-based approaches warrant further exploration. Our project is available at https://github.com/xiaomi/r1-aqa and https://huggingface.co/mispeech/r1-aqa.
@@ 1 Introduction
  • Based on test results from the MMAU test-mini benchmark, LALMs still lag far behind humans auditory-language reasoning, suggesting that the RL-based approaches warrant further exploration.
@@ 5 Conclusion
This report presents a case study on reinforcement learning and supervised fine-tuning for audio question answering. We directly apply the GRPO algorithm to Qwen2-Audio-7B-Instruct, and achieve state-of-the-art performance on the MMAU Test-mini benchmark, with an accuracy rate of 64.5%. This result indicates that reinforcement learning can be effectively applied to LALMs and audio multimodal, especially to AQA, which involves a significant generation-verification gap. Even if the model has a small number of parameters and the fine-tuning dataset is limited, it does not affect the implementation of reinforcement learning. In fact, on small-scale datasets, the strong generalization capability of reinforcement learning can be demonstrated even more clearly. However, the explicit reasoning process has not shown significant benefits for AQA tasks, and how to efficiently leverage deep thinking and step-by-step reasoning remains an open question for further research. LALMs still lag far behind humans auditory-language reasoning, suggesting that reinforcement learning warrants further exploration. Our future research will focus on the effective integration of the chain-of-thought paradigm into the audio modality.

### v2

@@ Abstract
Recently, reinforcement learning (RL) has been shown to greatly enhance the reasoning capabilities of large language models (LLMs), and RL-based approaches have been progressively applied to visual multimodal tasks. However, the audio modality has largely been overlooked in these developments. Thus, we conduct a series of RL explorations in audio understanding and reasoning, specifically focusing on the audio question answering (AQA) task. We leverage the group relative policy optimization (GRPO) algorithm to Qwen2-Audio-7B-Instruct, and our experiments demonstrated state-of-the-art performance on the MMAU Test-mini benchmark, achieving an accuracy rate of 64.5%. The main findings in this technical report are as follows: 1) The GRPO algorithm can be effectively applied to large audio language models (LALMs), even when the model has only 8.2B parameters; 2) With only 38k post-training samples, RL significantly outperforms supervised fine-tuning (SFT), indicating that RL-based approaches can be effective without large datasets; 3) The explicit reasoning process has not shown significant benefits for AQA tasks, and how to efficiently utilize deep thinking remains an open question for further research; 4) LALMs still lag far behind humans auditory-language reasoning, suggesting that the RL-based approaches warrant further explorations. Our project is available at https://github.com/xiaomi-research/r1-aqa and https://huggingface.co/mispeech/r1-aqa.
@@ 1 Introduction
  • Based on test results from the MMAU test-mini benchmark, LALMs still lag far behind humans auditory-language reasoning, suggesting that the RL-based approaches warrant further explorations.
@@ 5 Conclusion
This report presents a case study on reinforcement learning and supervised fine-tuning for audio question answering. We directly apply the GRPO algorithm to Qwen2-Audio-7B-Instruct, and achieve state-of-the-art performance on the MMAU Test-mini benchmark, with an accuracy rate of 64.5%. This result indicates that reinforcement learning can be effectively applied to LALMs and audio multimodal, especially to AQA, which involves a significant generation-verification gap. Even if the model has a small number of parameters and the fine-tuning dataset is limited, it does not affect the implementation of reinforcement learning. In fact, on small-scale datasets, the strong generalization capability of reinforcement learning can be demonstrated even more clearly. However, the explicit reasoning process has not shown significant benefits for AQA tasks, and how to efficiently leverage deep thinking and step-by-step reasoning remains an open question for further research. LALMs still lag far behind humans auditory-language reasoning, suggesting that reinforcement learning warrants further explorations. Our future research will focus on the effective integration of the chain-of-thought paradigm into the audio modality.